OpenStudy (anonymous):

The following system of equations is _____. y = -1/3x +2/3 2x + 6y = 4 independent equivalent inconsistent

OpenStudy (anonymous):

equivelent medal plz

OpenStudy (texaschic101):

y = -1/3x + 2/3 2x + 6y = 4 x + 3y = 2 3y = -x + 2 y = -1/3x + 2/3 These are equivalent equations :)